OpenStudy (anonymous):

A waiter earned $28 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 16 bills. 1 2 3 4

OpenStudy (anonymous):

@andijo76

OpenStudy (anonymous):

3 the answer is three

OpenStudy (anonymous):

sorry I am not andijo lol

OpenStudy (anonymous):

@andijo76 hey is my answer right?

OpenStudy (andijo76):

yes it is if you look at it this way 28=5x+1(16-x) 28=5x+16-1x 28-16=5x-1x 12=4x x=12/4 x=3

OpenStudy (anonymous):

thanks I should have explained

OpenStudy (andijo76):

its okay i just did so it can be used for future reference i always take notes so i know how to do similar problems

OpenStudy (anonymous):

okay thanks and thanks you for the medal

OpenStudy (andijo76):

your welcome

OpenStudy (anonymous):

lol